OpenStudy (anonymous):

If you pour a cup of coffee that is 200 F, and set it on desk in a room that is 68 F, and 10 minutes later it is 145 F, what temperature will it be 15 minutes after you originally poured it?

OpenStudy (anonymous):

if you lost 55 degrees in 10 min you would loose 5.5 degrees in 1 min then you multiply by 15. 15* 5.5 = 82.5 degrees then you subtract from 200 200-82.5 = 117.5 degrees in 15 min <--------Answer

zepdrix (zepdrix):

Newton's Law of Cooling:\[\Large\rm T(t)=T_a+(T_o-T_a)e^{-kt}\]Where: \(\Large\rm T_a\) is the surrounding or ambient temperature and \(\Large\rm T_o\) is the initial temperature of the thing in question (coffee in this case). For our problem we have:\[\Large\rm T_a=68\]\[\Large\rm T_o=200\] So this gives us an equation for temperature of:\[\Large\rm T(t)=68+(200-68)e^{-kt}\]\[\Large\rm T(t)=68+132e^{-kt}\]We're still missing our k coefficient though.

zepdrix (zepdrix):

They tell us that at time t=10 minutes, the temperature T=145 degrees. \[\Large\rm T(10)=145\]We'll use this to solve for our unknown k value.

zepdrix (zepdrix):

\[\Large\rm 145=68+132e^{-10k}\]So I plugged in 10 for t, and 145 for our T(t).

zepdrix (zepdrix):

Do you understand how to find your k value from here?

OpenStudy (crayla12):

k is .068